### v1

# AdS Length corrected Thermodynamics of a Black Hole in a Cloud of Strings and Perfect Fluid Dark Matter

Abstract
The research paper investigates the thermodynamic behavior of a black hole in a cloud of strings (CoS) and perfect fluid dark matter (PFDM), incorporating a minor correction to the Anti-de Sitter (AdS) length scale. The study reveals that while the entropy remains unaffected by the AdS length correction, other thermodynamic quantities such as enthalpy, free energy, pressure, internal energy, Gibbs free energy, and heat capacity exhibit significant modifications. Positive corrections to the AdS length ($\epsilon>0$) generally enhance these quantities, leading to faster growth or slower decay with increasing black hole size, while negative corrections ($\epsilon<0$) often produce the opposite effect, sometimes resulting in divergent or negative values. These findings highlight the sensitivity of black hole thermodynamics to geometric corrections and exotic matter fields, providing deeper information about the interplay between such modifications and thermodynamic stability. These results underscore the importance of considering minor corrections in theoretical models to better understand black hole behavior in the presence of dark matter and string-like matter distributions.
Keywords: Cloud of Strings (CoS), Perfect Fluid Dark Matter (PFDM), AdS length correction, Quantum/ Thermal Corrections, Black Hole Thermodynamics.
1 Introduction
The study of black hole thermodynamics [1] in various modified gravity and matter field backgrounds has gained significant momentum in recent decades. This interest is driven both by the theoretical quest to unify general relativity with quantum field theory [2]-[10] and by astrophysical observations indicating the presence of dark matter and other exotic energy distributions in the universe.
Among the various matter models considered in the context of black hole spacetimes, the cloud of strings (CoS) and perfect fluid dark matter (PFDM) have emerged as compelling candidates. A cloud of strings represents a classical limit of string theory, offering a simplified model for string-like matter distributions that can arise in various high-energy scenarios [11]-[17]. On the other hand, PFDM has been proposed as a phenomenological model that effectively captures the gravitational influence of dark matter at both galactic and cosmological scales [18]-[24].
Incorporating such matter sources into black hole spacetimes leads to intriguing modifications in both geometric and thermodynamic properties. The presence of CoS and PFDM alters the black hole metric functions, influencing quantities like the horizon radius, surface gravity, and ultimately the Hawking temperature [25] and entropy [26].
Furthermore, within the framework of extended black hole thermodynamics, where the cosmological constant $\Lambda$ is treated as a thermodynamic pressure [27]-[31], the Anti-de Sitter (AdS) length scale $l$ plays a central role. Motivated by quantum corrections or higher-order effects, it is reasonable to consider minor modifications to the AdS length. These corrections, albeit small, can have significant implications for the thermodynamic behavior of the black hole, potentially affecting stability criteria, phase transitions, and critical phenomena [32]-[38]. Moreover, the minor corrections in AdS length have led to the universal relations [39].
In this paper, we investigate a static, spherically symmetric black hole solution surrounded by both a cloud of strings and perfect fluid dark matter [40]. We introduce a perturbative correction to the AdS length scale and study its impact on the thermodynamics of the system. Our goal is to analyze how such a correction influences the temperature, entropy, and heat capacity of the black hole, thereby shedding light on the interplay between the background matter fields and geometric corrections in AdS space.
The structure of the paper is as follows: In Section 2, we briefly review the black hole solution with CoS and PFDM in an AdS background. In Section 3, we incorporate a small correction to the AdS length and derive and analyze the modified thermodynamic quantities. We also analyze stability of the BH. We present our summarized results and outlook in Section 4.
2 Black Hole with Cloud of Strings and PFDM
The line element describing an AdS charged black hole (BH) in the presence of a cloud of strings (CoS) and perfect fluid dark matter (PFDM), as presented in [40], takes the following form:
$ds^{2}=$ $-\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}-a+\frac{\lambda}{r}\ln\left(\frac{r}{\lambda}\right)+\frac{r^{2}}{l^{2}}\right)dt^{2}$
$+\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}-a+\frac{\lambda}{r}\ln\left(\frac{r}{\lambda}\right)+\frac{r^{2}}{l^{2}}\right)^{-1}dr^{2}+r^{2}d\Omega^{2}$
Here, $a$ represents the CoS parameter, $l$ is the AdS radius, $Q$ denotes the electric charge, and $\lambda$ characterizes the PFDM scale parameter.
Solving $f(r)=0$ at the event horizon radius $r_{+}$ yields the mass of the black hole:
$M_{+}=\frac{r_{+}}{2}\left(1+\frac{Q^{2}}{r_{+}^{2}}-a+\frac{r_{+}^{2}}{l^{2}}+\frac{\lambda}{2r_{+}}\ln\left(\frac{r_{+}}{\lambda}\right)\right)$ (1)
By evaluating the surface gravity at the horizon, the Hawking temperature is derived using $T_{H}=\frac{f^{\prime}(r_{+})}{4\pi}$:
$T_{H}=\frac{1}{4\pi r_{+}}\left(1-a+\frac{3r_{+}^{2}}{l^{2}}-\frac{Q^{2}}{r_{+}^{2}}+\frac{\lambda}{r_{+}}\right)$ (2)
Incorporating the first law of thermodynamics for this system:
$dM_{+}=T_{+}dS_{+}+\Phi dQ+\lambda d\Lambda_{1}+PdV,$
where $\Lambda_{1}$ is conjugate to $\lambda$, and $\Phi$ is the electric potential.
Assuming constant charge $Q$, the black hole entropy becomes:
$S_{0}=\int\frac{1}{T_{+}}\frac{dM_{+}}{dr_{+}}dr_{+}=\pi r_{+}^{2}=\frac{A_{+}}{4}$ (3)
Integrating the first law leads to the expression for enthalpy:
$H_{0}=\int T_{H}\,dS_{0}=\frac{1}{2}\left(\lambda\ln r_{+}+\frac{Q^{2}}{r_{+}}+(1-a)r_{+}+\frac{r_{+}^{3}}{l^{2}}\right)$ (4)
The Helmholtz free energy, obtained via a Legendre transform, is given by:
$F_{0}=-\int S_{0}\,dT_{H}=\frac{1}{4}\left(2\lambda\ln r_{+}+\frac{3Q^{2}}{r_{+}}+(1-a)r_{+}-\frac{r_{+}^{3}}{l^{2}}\right)$ (5)
The thermodynamic volume of the black hole is:
$V=\frac{4}{3}\pi r_{+}^{3}$ (6)
Using the relation $P_{0}=-\left.\frac{dF_{0}}{dV}\right|_{T_{H}}$, the thermodynamic pressure is found to be:
$P_{0}=-\frac{1-a-\frac{3Q^{2}}{r_{+}^{2}}+\frac{2\lambda}{r_{+}}-\frac{3r_{+}^{2}}{l^{2}}}{16\pi r_{+}^{2}}$ (7)
The internal energy is computed from $U_{0}=H_{0}-P_{0}V$:
$U_{0}=\frac{1}{12}\left(2\lambda(1+3\ln r_{+})+\frac{3Q^{2}}{r_{+}}-7(a-1)r_{+}+\frac{3r_{+}^{3}}{l^{2}}\right)$ (8)
The Gibbs free energy is determined using $G_{0}=F_{0}+P_{0}V$:
$G_{0}=$ $-\frac{1}{12}r_{+}\left(1-a-\frac{3Q^{2}}{r_{+}^{2}}+\frac{2\lambda}{r_{+}}-\frac{3r_{+}^{2}}{l^{2}}\right)$
$+\frac{1}{4}\left(2\lambda\ln r_{+}+\frac{3Q^{2}}{r_{+}}+(1-a)r_{+}-\frac{r_{+}^{3}}{l^{2}}\right)$ (9)
Lastly, the heat capacity $C_{0}=T_{H}\left(\frac{dS}{dT_{H}}\right)$, which informs us about the thermodynamic stability of the black hole, is given as:
$C_{0}=\frac{2\pi r_{+}^{2}\left(-l^{2}Q^{2}+l^{2}\lambda r_{+}+(1-a)l^{2}r_{+}^{2}+3r_{+}^{4}\right)}{3l^{2}Q^{2}-2l^{2}\lambda r_{+}+(a-1)l^{2}r_{+}^{2}+3r_{+}^{4}}$ (10)
3 AdS length Corrected Thermodynamics
Replacing AdS length by effective AdS length $l\to l/\sqrt{1+\epsilon}$, where $\epsilon\ll 1$ is the correction parameter, we get the BH solution surrounded by CoS and PFDM
$\begin{split}ds^{2}=&-\left[1-a-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}+\frac{\lambda}{r}\ln\left(\frac{r}{\lambda}\right)+\frac{r^{2}}{l^{2}}(1+\epsilon)\right]dt^{2}\\ &+\left[1-a-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}+\frac{\lambda}{r}\ln\left(\frac{r}{\lambda}\right)+\frac{r^{2}}{l^{2}}(1+\epsilon)\right]dr^{2}+r^{2}d\Omega^{2}\end{split}$ (11)
Using $f(r_{+})=0$, we get the BH mass
$M=\frac{(1-a)}{2}r_{+}+\frac{Q^{2}}{2r_{+}}+\frac{\lambda}{2}\ln\left(\frac{r_{+}}{\lambda}\right)+\frac{r_{+}^{3}}{2l^{2}}(1+\epsilon)$ (12)
The equation $\left.\frac{f^{\prime}(r)}{4\pi}\right|_{r=r_{+}}$ gives the corrected Hawking Temperature
$T_{H}=\frac{1-a}{4\pi r_{+}}-\frac{Q^{2}}{4\pi r_{+}^{3}}+\frac{\lambda}{4\pi r_{+}^{2}}+\frac{3r_{+}}{4\pi l^{2}}(1+\epsilon)$ (13)
At constant charge, the entropy of the black hole with correction to AdS length is given by:
$S_{c}=\int\frac{1}{T_{H}}\frac{dM}{dr_{+}}dr_{+}$ (14)
Using (13) and (12) and integrating, we get the corrected entropy
$S_{c}=\pi r_{+}^{2}$ (15)
Here we find that the correction to the AdS length does not influence the entropy.
The corrected enthalpy is mathematically given by:
$H_{c}=\int T_{H}\,dS_{c}$ (16)
Evaluating the integral together with (13) and (15) gives:
$H_{c}=\frac{1}{2}\left[\lambda\ln\left(r_{+}\right)+\frac{Q^{2}}{r_{+}}+r_{+}-ar_{+}+\frac{(1+\epsilon)r_{+}^{3}}{l^{2}}\right]$ (17)
To observe the effect of a minor correction to the AdS length on the black hole enthalpy, we plot a graph of (17) versus the horizon radius for different values of the correction parameter $\epsilon$.
Figure 1: Enthalpy $H_{c}$ v/s horizon radius $r_{+}$ with different values of $\epsilon$. The curves show the behavior for $\epsilon=0$ (uncorrected), $\epsilon>0$, and $\epsilon<0$.
From the graph (1), we observe that the three curves converge for small black holes, exhibiting positive asymptotic behavior and decrease with horizon radius. While the $\epsilon$ curve continues decreasing and becomes more negative with increasing black hole size, the $\epsilon$ curve demonstrates the same trend as the uncorrected one, increasing from $r_{+}\approx 0.5$, growing faster and becoming more positive with increasing black hole size.
The corrected free energy is mathematically given by:
$F_{c}=-\int S_{c}\,dT_{H}$ (18)
Evaluating the integral together with (13) and (15) gives:
$F_{c}=-\frac{1}{4}\left[2\lambda\ln\left(r_{+}\right)+\frac{3Q^{2}}{r_{+}}+r_{+}-ar_{+}-\frac{(1+\epsilon)r_{+}^{3}}{l^{2}}\right]$ (19)
To observe the effect of a minor correction to the AdS length on the black hole free energy, we plot a graph of (19) versus the horizon radius for different values of the correction parameter $\epsilon$.
Figure 2: Free energy $F_{c}$ v/s horizon radius $r_{+}$ with different values of $\epsilon$. The curves show the behavior for $\epsilon=0$ (uncorrected), $\epsilon>0$, and $\epsilon<0$.
From the graph (2), we observe that the three curves converge for small black holes, exhibiting positive asymptotic behavior and decrease with horizon radius. While the $\epsilon$ curve demonstrates the same trend as the uncorrected one, continuing to decrease and becoming more negative with increasing black hole size, dropping faster, the $\epsilon$ curve shows the opposite trend by increasing from $r_{+}\approx 0.5$ and becomes more positive with increasing black hole size.
The volume is given by:
$V=\frac{4}{3}\pi r_{+}^{3}$ (20)
The corrected pressure is given by:
$P_{c}=-\left.\frac{dF_{c}}{dV}\right|_{T}$ (21)
Using the expression for volume together with (19), we get:
$P_{c}=-\frac{1}{16\pi r_{+}^{2}}\left[1-a-\frac{3Q^{2}}{r_{+}^{2}}+\frac{2\lambda}{r_{+}}-\frac{3(1+\epsilon)r_{+}^{2}}{l^{2}}\right]$ (22)
To observe the effect of a minor correction to the AdS length on the black hole pressure, we plot a graph of (22) versus the horizon radius for different values of the correction parameter $\epsilon$.
Figure 3: Pressure $P_{c}$ v/s the horizon radius $r_{+}$ with different values of $\epsilon$. The curves show the behavior for $\epsilon=0$ (uncorrected), $\epsilon>0$, and $\epsilon<0$.
From the graph (3), we observe that all three curves show positive asymptotic behavior for small black holes, decreasing with horizon radius and almost attain their respective constant values beyond $r_{+}\approx 0.7$. As the black hole size grows, the positive correction keeps the curve above the uncorrected one, thereby making it drop slower with increasing size, while the negative correction decreases the pressure more and even makes it go into negative values with increasing size.
The corrected internal energy is given by:
$U_{c}=H_{c}-P_{c}V$ (23)
Using (17) and (22) together with the expression for volume, we get:
$U_{c}=\frac{1}{12}\left[2\lambda(1+3\ln\left(r_{+}\right))+\frac{3Q^{2}}{r_{+}}-7(-1+a)r_{+}+\frac{3(1+\epsilon)r_{+}^{3}}{l^{2}}\right]$ (24)
To observe the effect of a minor correction to the AdS length on the black hole internal energy, we plot a graph of (24) versus the horizon radius for different values of the correction parameter $\epsilon$.
Figure 4: Internal energy $U_{c}$ v/s horizon radius $r_{+}$ with different values of $\epsilon$. The curves show the behavior for $\epsilon=0$ (uncorrected), $\epsilon>0$, and $\epsilon<0$.
From the graph (4), we observe that the three curves converge for small black holes, exhibiting positive asymptotic behavior and decrease with horizon radius. While the $\epsilon$ curve continues decreasing and becomes more negative with increasing black hole size, the $\epsilon$ curve demonstrates the same trend as the uncorrected one, increasing from $r_{+}\approx 0.5$, growing faster and becoming more positive with increasing black hole size.
The corrected Gibbs free energy is evaluated as:
$G_{c}=H_{c}-T_{H}S_{c}$ (25)
Inserting values, we get:
$G_{c}=\frac{1}{4}\left[\lambda(-1+2\ln(r_{+}))+\frac{3Q^{2}}{r_{+}}-(-1+a)r_{+}-\frac{(1+\epsilon)r_{+}^{3}}{l^{2}}\right]$ (26)
To observe the effect of a minor correction to the AdS length on the black hole Gibbs free energy, we plot a graph of (26) versus the horizon radius for different values of the correction parameter $\epsilon$.
Figure 5: Gibbs free energy $G_{c}$ v/s horizon radius $r_{+}$ with different values of $\epsilon$. The curves show the behavior for $\epsilon=0$ (uncorrected), $\epsilon>0$, and $\epsilon<0$.
From the graph (5), we observe that the three curves converge for small black holes, exhibiting positive asymptotic behavior and decrease with horizon radius. While the $\epsilon$ curve demonstrates the same trend as the uncorrected one, continuing to decrease and becoming more negative with increasing black hole size, dropping faster, the $\epsilon$ curve shows the opposite trend by increasing from $r_{+}\approx 0.5$ and becomes more positive with increasing black hole size.
The expression for the corrected heat capacity is:
$C_{c}=T_{H}\frac{dS_{c}}{dT_{H}}=T_{H}\frac{\left(dS_{c}/dr_{+}\right)}{\left(dT_{H}/dr_{+}\right)}$ (27)
Inserting the values, we get:
$C_{c}=\frac{2(-l^{2}Q^{2}+l^{2}\lambda r_{+}-(-1+a)l^{2}r_{+}^{2}+3r_{+}^{4})(-\gamma+\pi\beta r_{+}^{2}+\pi^{2}r_{+}^{4})}{\pi r_{+}^{2}(3l^{2}Q^{2}-2l^{2}\lambda r_{+}+(-1+a)l^{2}r_{+}^{2}+3r_{+}^{4})}$ (28)
To observe the effect of a minor correction to the AdS length on the black hole heat capacity, we plot a graph of (28) versus the horizon radius for different values of the correction parameter $\epsilon$.
Figure 6: Heat capacity $C_{c}$ v/s of horizon radius $r_{+}$ with different values of $\epsilon$. The curves show the behavior for $\epsilon=0$ (uncorrected), $\epsilon>0$, and $\epsilon<0$.
From the graph (6), we observe that the three curves converge for small horizon radii, being negligible, and depart from one another at $r_{+}\approx 0.3$. The curve corresponding to $\epsilon$ shows the same trend as the uncorrected curve but increases faster and becomes more positive with increasing black hole size and converges with it at large horizon radii. The negative correction exhibits divergence at $r_{+}\approx 0.75$ and gets closer to the other curves at large size.
4 Conclusions
The research paper investigated the thermodynamic behavior of a black hole in by a cloud of strings (CoS) and perfect fluid dark matter (PFDM), incorporating a minor correction to the Anti-de Sitter (AdS) length scale. The study revealed that while the entropy remains unaffected by the AdS length correction, other thermodynamic quantities such as enthalpy, free energy, pressure, internal energy, Gibbs free energy, and heat capacity exhibited significant modifications. Positive corrections to the AdS length ($\epsilon>0$) generally enhanced these quantities, leading to faster growth or slower decay with increasing black hole size, while negative corrections ($\epsilon<0$) often produced the opposite effect, sometimes resulting in divergent or negative values. These findings highlight the sensitivity of black hole thermodynamics to geometric corrections and exotic matter fields, providing deep information about the interplay between such modifications and thermodynamic stability. The results underscore the importance of considering minor corrections in theoretical models to better understand black hole behavior in the presence of dark matter and string-like matter distributions.
Acknowledgment
N.A.G. is grateful to wife Mymoona for editing and proof-reading the manuscript.